# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v3

@@ Abstract
Autonomous mobile robots must maintain safety, but should not sacrifice performance, leading to the classical reach-avoid problem. This paper seeks to compute trajectory plans for which a robot is guaranteed to reach a goal and to avoid obstacles in the specific near danger case, also known as a narrow gap, where the agent starts near the goal, but must navigate through tight obstacles that block its path. The proposed method builds off of a common approach of using a simplified planning model to generate plans, which are then tracked using a high-fidelity tracking model and controller. Existing safe planning approaches use reachability analysis to overapproximate the error between these models, but this introduces additional numerical approximation error and thereby conservativeness that prevents goal reaching. The present work instead proposes a Piecewise Affine Reach-avoid Computation (PARC) method to tightly approximate the reachable set of the planning model. PARC significantly reduces conservativeness through a careful choice of the planning model and set representation, along with an effective approach to handling time-varying tracking errors. The utility of this method is demonstrated through extensive numerical experiments in which PARC outperforms state-of-the-art reach avoid methods in near-danger goal reaching. Furthermore, in a simulated demonstration, PARC enables generation of provably-safe extreme vehicle dynamics drift parking maneuvers. A preliminary hardware demo on a TurtleBot3 also validates the method.
@@ I Introduction
The increasing deployment of autonomous robots in daily tasks necessitates addressing complex operational challenges. Many of these tasks can be conceptualized as reach-avoid problems, wherein a robot must steer its states into a set of goal states without violating safety-critical constraints [fisac2015reach]. Achieving both liveness – the assurance of goal fulfillment [alpern1985defining] – and adherence to safety constraints is pivotal in these scenarios. In this paper, we are particularly interested in collision-free goal reaching near danger, meaning that the agent starts near the goal, but must navigate through tight obstacles—typically within the width of a robot’s body—that block its path. As we will show, this problem setup is very difficult for existing robust planning and control approaches to provide reach-avoid guarantees on general dynamical systems due to over-conservativeness. We note that the specific “near danger” condition where these planners and controllers fail is task-dependent. Some examples include flying a quadrotor drone through a narrow gap, or parallel parking an autonomous car.
@@ I-A Reach-Avoid Approaches and Challenges
Figure 1: Our overall goal is to find safe goal-reaching motion plans near danger. In (a), we show the components of our proposed Piecewise Affine Reach-avoid Computation (PARC) method, with the paper section number on the top left, and the corresponding symbol on the bottom. This method enables a new state-of-the-art for safe reach-avoid problems, as evidenced by (b), which shows examples of extreme vehicle maneuvers. These planned drift parking trajectories are sampled from the Backward Reach-Avoid Set (BRAS, green triangle on the top left) using PARC. Our vehicle is in a blue time lapse. The two red cars are obstacles. The goal set for the center of mass is the green rectangle. The small size of the BRAS indicates the challenge of finding safe trajectories.
@@
Both control intervention and planner-tracker approaches suffer from challenges in numerical representation due to the need to underapproximate the set of states that reach the target set and overapproximate those that avoids obstacles when guaranteeing safety and liveness in implementation. Implementations such as sums-of-squares polynomials [majumdar2017funnel, kousik2020bridging, landry2018reach], grid-based PDE solutions [chen2021fastrack], or polytope methods [kousik2019safe, althoff2021set] all introduce additional conservativeness purely due to compensating for numerical representation error. This challenge is exacerbated for systems with high-dimensional, nonlinear dynamics or observations; machine learning can address some of these concerns [bansal2021deepreach, michaux2023reachability, dawson2022safe, xiao2023barriernet, selim2022safe], at the expense of losing guarantees.
@@ I-B Proposed Method Summary and Context
Our method fits in the context of both control and planning. We build off of a long history of discrete-time PWA system analysis in optimal control (e.g., [thomas2006robust, kerrigan2002optimal, rakovic2005robust]), which we extend to the continuous-time reach-avoid problem for robot motion planning. Furthermore, our approach is complementary to sampling-based motion planners such as RRT [lavalle1998rapidly]; our approach could be applied in the future to compute a safety funnel around a sampled motion plan, or we could leverage sampling to seed our reachability analysis.
@@ I-C Contributions
    We stress-test PARC on examples drawn from the safe control and planning literature: a 4-D TurtleBot, a 6-D planar quadrotor, a 10-D near-hover quadrotor, and a 13-D general quadrotor. We show that PARC outperforms state-of-the-art methods such as FaSTrack [chen2021fastrack], Neural CLBF [dawson2022safe], RTD [kousik2020bridging, kousik2019safe], and KDF [verginis2022kdf] in near danger reach-avoid tasks due to reduced conservativeness from the combination of planning model and set representation choice and better implementation of tracking error. We also demonstrate PARC on a simulated 6-D vehicle model to plan safe, extreme drift parallel parking maneuvers, and on a 4-D physical TurtleBot as a preliminary hardware experiment.
Paper Organization: Next, we introduce mathematical objects used throughout our formulation (Section II). We then present our problem formulation (Section III) and proposed approach (Section IV). We assess PARC with numerical experiments (Section V), then demonstrate a drift parking extreme vehicle dynamics example in simulation, extending beyond the state of the art in reachability-based safe motion-planning and control methods (Section VI). We also showcase PARC’s ability to extend to real robots with a preliminary hardware experiment (Section VII). Finally, we discuss takeaways and limitations in Section VIII. Note, we have extensive appendices, and call out key results in the text.
@@ II Preliminaries
We now introduce our notation conventions and define H-polytopes and PWA functions, which we use to represent sets and dynamics of planning model throughout the paper. We then present a one-step Backward Reachable Set (BRS) computation that is the core of our proposed PARC method.
@@ II-C PWA Systems
We consider discrete-time, time-variant PWA functions to enable our reach-avoid analysis. We denote a PWA function as $\psi:X_{0}\to X_{t_{\mathrm{\textnormal{f}}}}\subseteq\mathbb{R}^{{n}}$ with output $\mathbf{x}(t_{\mathrm{\textnormal{f}}})=\mathbf{\psi}\!\left(\mathbf{x}(0)\right)$ given an input $\mathbf{x}(0)=\mathbf{x}_{0}\in X_{0}\subset\mathbb{R}^{{n}}$. The system has state $\mathbf{x}(t)\in X_{t}\subset\mathbb{R}^{{n}}$ at time $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, which is the solution of difference equation of dynamics:
@@
where $\mathbf{C}_{\ell,t}\in\mathbb{R}^{{n}\times{n}}$ and $\mathbf{d}_{\ell,t}\in\mathbb{R}^{{n}}$ define the dynamics of the $\ell^{\mathrm{\textnormal{th}}}$ PWA region. The index $\ell$ is given by
@@
where $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$ is the $i^{\mathrm{\textnormal{th}}}$ of the ${n}_{\mathrm{\textnormal{PWA}},t}$ PWA regions at time $t$. We say that $\mathbf{x}(t)$ has the mode $s_{t}=\ell$ at time $t$.
In addition, if the mode of $\mathbf{x}(0),\mathbf{x}(\Delta t),\cdots,\mathbf{x}(t_{\mathrm{\textnormal{f}}}-\Delta t)$ is $s_{0},s_{\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t}$ for some input $\mathbf{x}(0)\in X_{0}$, we denote the mode sequence for $\mathbf{x}(0)$ as:
@@
The reduction of a PWA system to a specific affine dynamics is the core idea that enables the efficient underapproximating BRS computation of PARC. Note, taking advantage of fixed mode sequences is well-established in robotics and controls (e.g., [thomas2006robust, desimini2020robust, kerrigan2002optimal, wensing2023optimization]), but we believe we are the first to show its utility in safe reach-avoid computation of robot planning models.
We require that the interior of each PWA region does not intersect with other regions, such that any two regions are either disjoint or intersect only at their boundaries:
$\left\{\mathbf{x}\ |\ \mathbf{A}_{i}(t)\mathbf{x}<\mathbf{b}_{i}(t)\right\}\cap\mathcal{P}(\mathbf{A}_{j}(t),\mathbf{b}_{j}(t))=\emptyset,$ (6)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ and $i,j\in\{1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\}$, $i\neq j$. Thus, the $\min$ in (3) dictates the dynamics’ behavior at the intersecting subsets of boundaries.
Note that, for $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, we define:
$X_{t}=\bigcup_{i=1}^{{n}_{\mathrm{\textnormal{PWA}},t}}\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t}),$ (7)
such that the input to $\psi$ is restricted to $\{\mathbf{x}_{0}\ |\ \mathbf{x}_{t}\in X_{t},t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t\}$. Thus, a PWA function $\psi$ can be fully defined and parameterized by some given $\mathbf{C}_{i,t}$, $\mathbf{d}_{i,t}$, $\mathbf{A}_{i,t}$, and $\mathbf{b}_{i,t}$ for each $i=1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}$ and $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$.
As a particular case of this problem formulation, when ${n}_{\mathrm{\textnormal{PWA}},t}=1$ for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$, one obtains the discrete-time, time-variant affine system (LABEL:eq:time_variant_affine) from the PWA system (2) and (3) automatically, as $S=(1,\cdots,1)$ always.
@@
As will be shown in Section IV-C, the use of linear interpolation to bridge between discrete and continuous time is necessary to enable avoid set computation with H-polytopes. Similar reachability techniques to ours have been shown to work with Bernstein polynomials [michaux2023can], but only for forward reachability, to the best of our knowledge. Nonetheless, it may still be possible for PARC to handle alternative interpolation methods such as cubic splines or Bézier curves by using other set representations such as polynomial zonotopes [kochdumper2020sparse]. In this paper, we opted for the most straightforward approach of linear interpolation and leave the investigation with other interpolation methods as future work.
@@ II-D One Step BRS via Inverse Affine Map
Consider an affine system $\mathbf{x}(t+\Delta t)=\mathbf{C}\mathbf{x}(t)+\mathbf{d}$. Define the one-step BRS as the map $\mathcal{B}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{C},\mathbf{d}\right)=\left\{\mathbf{x}|\ \mathbf{A}\mathbf{x}_{next}\leq\mathbf{b},\ \mathbf{x}_{next}=\mathbf{C}\mathbf{x}+\mathbf{d}\right\}$. This one-step BRS is in fact an H-polytope:
@@ III-A Planning Model
Changing the workspace state either does not affect the planning dynamics $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ in the workspace or else the dynamics are constant:
$\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\mathbf{f}_{\mathbf{w}}\right\|_{2}=0\quad\mathrm{\textnormal{and}}\quad\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{p}}\right\|_{\mathrm{\textnormal{F}}}=0,$ (13)
@@
This is not overly restrictive; a variety of ETI planning models are shown in Section V. The need for the dynamics to be constant in some states is because we treat $\mathbf{k}$ as states with 00 dynamics to formulate our reachability method.
@@ III-B Tracking Model
where $\mathbf{z}_{0}\in Z\subset\mathbb{R}^{{n}_{z}}$ is the initial state. The tracking state $\mathbf{z}$ includes the planning state $\mathbf{p}$, i.e., $\mathbf{z}=[\mathbf{p}^{\intercal},\mathbf{z}_{\mathrm{\textnormal{other}}}^{\intercal}]^{\intercal}$. Given a plan $\mathbf{p}$ parameterized by $\mathbf{k}$, we use a feedback tracking controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}(t,\mathbf{z}(t);\mathbf{k})\in U$ (e.g., PID or MPC) to control the tracking model. We assume $Z$ and $U$ are compact and the dynamics are Lipschitz such that trajectories exist [grant2014theory]. For a particular $\mathbf{z}_{0}$, we denote the resulting, unique realized trajectory (i.e., solution to (14)) as $\mathbf{z}(t;\mathbf{z}_{0},\mathbf{k})$ at time $t$.

### v4

@@ Abstract
Autonomous mobile robots must maintain safety, but should not sacrifice performance, leading to the classical reach-avoid problem: find a trajectory that is guaranteed to reach a goal and avoid obstacles. This paper addresses the near danger case, also known as a narrow gap, where the agent starts near the goal, but must navigate through tight obstacles that block its path. The proposed method builds off the common approach of using a simplified planning model to generate plans, which are then tracked using a high-fidelity tracking model and controller. Existing approaches use reachability analysis to overapproximate the error between these models and ensure safety, but doing so introduces numerical approximation error conservativeness that prevents goal-reaching. The present work instead proposes a Piecewise Affine Reach-avoid Computation (PARC) method to tightly approximate the reachable set of the planning model. PARC significantly reduces conservativeness through a careful choice of the planning model and set representation, along with an effective approach to handling time-varying tracking errors. The utility of this method is demonstrated through extensive numerical experiments in which PARC outperforms state-of-the-art reach avoid methods in near-danger goal reaching. Furthermore, in a simulated demonstration, PARC enables the generation of provably-safe extreme vehicle dynamics drift parking maneuvers. A preliminary hardware demo on a TurtleBot3 also validates the method.
@@ I Introduction
The increasing deployment of autonomous robots in daily tasks necessitates addressing complex operational challenges. Many of these tasks can be conceptualized as reach-avoid problems, wherein a robot must steer its states into a set of goal states without violating safety-critical constraints [fisac2015reach]. Achieving both liveness—the assurance of goal fulfillment [alpern1985defining]—and adherence to safety constraints is pivotal in these scenarios. In this paper, we are particularly interested in collision-free goal reaching near danger, meaning that the agent starts near the goal, but must navigate past tightly-spaced obstacles, typically less than the width of a robot’s body from the goal. We assume an alternative approach (e.g., [kousik2019safe, kousik2020bridging, chen2021fastrack]) has safely brought the robot near the goal but cannot complete the final goal-reaching maneuver. As we will show, this problem setup is very difficult for existing robust planning and control approaches to provide reach-avoid guarantees on general dynamical systems due to over-conservativeness. We note that the specific “near danger” condition where these planners and controllers fail is task-dependent. Some examples include flying a quadrotor drone through a narrow gap, or parallel parking an autonomous car.
@@ I-A Reach-Avoid Approaches and Challenges
Figure 1: Our overall goal is to find safe goal-reaching motion plans near danger. In (a), we show the components of our proposed Piecewise Affine Reach-avoid Computation (PARC) method, with the paper section number on the top left, and the corresponding symbol on the bottom. This method enables a new state-of-the-art for safe reach-avoid problems, as evidenced by (b), which shows examples of extreme vehicle maneuvers. These planned drift parking trajectories are sampled from the Backward Reach-Avoid Set (BRAS, green triangle on the top left) using PARC. Our vehicle is shown with a blue timelapse. The two red cars are obstacles. The goal set for the center of mass is the green rectangle. The small size of the BRAS indicates the challenge of finding safe trajectories.
@@
Both control intervention and planner-tracker approaches suffer from challenges in numerical representation due to the need to underapproximate the set of states that reach the target set and overapproximate those that avoids obstacles when guaranteeing safety and liveness in implementation. Implementations such as sums-of-squares polynomials [majumdar2017funnel, kousik2020bridging, landry2018reach], grid-based partial differential equation (PDE) solutions [chen2021fastrack], or set-based methods [kousik2019safe, althoff2021set] all introduce additional conservativeness purely due to compensating for numerical representation error. This challenge is exacerbated for systems with high-dimensional, nonlinear dynamics or observations; machine learning can address some of these concerns [bansal2021deepreach, michaux2023reachability, dawson2022safe, xiao2023barriernet, selim2022safe], at the expense of losing guarantees.
@@ I-B Proposed Method Summary and Context
Our method fits in the context of both control and planning. We build off of a long history of discrete-time PWA system analysis in optimal control (e.g., [thomas2006robust, kerrigan2002optimal, rakovic2005robust]), in which we extend to the continuous-time reach-avoid problem for robot motion planning. Furthermore, our approach is complementary to sampling-based motion planners such as rapidly exploring random tree (RRT) [lavalle1998rapidly]; our approach could be applied in the future to compute a safety funnel around a sampled motion plan, or we could leverage sampling to seed our reachability analysis.
@@ I-C Contributions
    We stress-test PARC on examples drawn from the safe control and planning literature: a 4-D TurtleBot, a 6-D planar quadrotor, a 10-D near-hover quadrotor, and a 13-D general quadrotor. We show that PARC outperforms state-of-the-art methods such as Fast and Safe Tracking (FaSTrack) [chen2021fastrack], Neural Control Lyapunov Barrier Function (Neural CLBF) [dawson2022safe], Reachability-based Trajectory Design (RTD) [kousik2020bridging, kousik2019safe], and KinoDynamic motion planning via Funnel control (KDF) [verginis2022kdf] in near danger reach-avoid tasks due to reduced conservativeness from the combination of planning model and set representation choice and better implementation of tracking error.
  4. 4.
    We demonstrate PARC on a simulated 6-D vehicle model to plan safe, extreme drift parallel parking maneuvers, and on a 4-D physical TurtleBot as a preliminary hardware experiment.
Paper Organization: Next, we introduce mathematical objects used throughout our formulation (Section II). We then present our problem formulation (Section III) and proposed approach (Section IV). We assess PARC with numerical experiments (Section V), then demonstrate a drift parking extreme vehicle dynamics example in simulation, extending beyond the state of the art in reachability-based safe motion-planning and control methods (Section VI). We also showcase PARC’s ability to extend to real robots with a preliminary hardware experiment (Section VII). Finally, we discuss takeaways and limitations in Section VIII. For presentation’s sake, we move most proofs and experiment details to appendices, and call out key results in the text.
@@ II Preliminaries
We now introduce our notation conventions and define H-polytopes and PWA system, which we use to represent sets and dynamics of planning model throughout the paper. We then present a one-step Backward Reachable Set (BRS) computation that is the core of our proposed PARC method.
@@ II-C PWA Systems
In this section, we formally define the discrete-time, time-variant PWA system, which becomes the main objective of our reachability analysis.
The state of the system, $\mathbf{x}(t)\in X\subset\mathbb{R}^{{n}}$, evolves at discrete times $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ according to the difference equation:
@@
where $\mathbf{C}_{\ell,t}\in\mathbb{R}^{{n}\times{n}}$ and $\mathbf{d}_{\ell,t}\in\mathbb{R}^{{n}}$ define the dynamics of the $\ell^{\mathrm{\textnormal{th}}}$ PWA region at time $t$. The index $\ell$ is given by
@@
where $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$ is the $i^{\mathrm{\textnormal{th}}}$ of the ${n}_{\mathrm{\textnormal{PWA}},t}$ PWA regions at time $t$. We ensure that the PWA regions form a tessellation of the domain $X$ at each time step $t$, meaning their union covers the entire domain and their interiors do not intersect:
$X=\bigcup_{i=1}^{{n}_{\mathrm{\textnormal{PWA}},t}}$ $\mathcal{P}(\mathbf{A}_{i,t},\mathbf{b}_{i,t})$ (4)
$\left\{\mathbf{x}\ |\ \mathbf{A}_{i,t}\mathbf{x}<\mathbf{b}_{i,t}\right\}\cap$ $\mathcal{P}(\mathbf{A}_{j,t},\mathbf{b}_{j,t})=\emptyset,$ (5)
for all $t=0,\Delta t,\cdots,t_{\mathrm{\textnormal{f}}}-\Delta t$ and $i,j\in\{1,\dots,{n}_{\mathrm{\textnormal{PWA}},t}\}$, $i\neq j$. With this condition and the decision rule (3), the PWA-system has well-defined dynamics on the domain $X$. Hence, for a given initial state $\mathbf{x}_{0}$, the PWA system has a unique state at each time $t$, which we denote as $\mathbf{x}(t;\mathbf{x}_{0})$. We drop $\mathbf{x}_{0}$ and simplify it to $\mathbf{x}(t)$ when it is contextually evident.
We say that the state $\mathbf{x}(t)$ has the mode $s_{t}=\ell$ if $\mathbf{x}(t)\in\mathcal{P}(\mathbf{A}_{\ell,t},\mathbf{b}_{\ell,t})$. Further, if the mode of $\mathbf{x}(0),\mathbf{x}(\Delta t),\cdots,\mathbf{x}(t_{\mathrm{\textnormal{f}}}-\Delta t)$ is $s_{0},s_{\Delta t},\cdots,s_{t_{\mathrm{\textnormal{f}}}-\Delta t}$ for some input $\mathbf{x}(0)\in X$, we denote the mode sequence for $\mathbf{x}(0)$ as:
@@
The reduction of a PWA system to specific affine dynamics is the core idea that enables the efficient underapproximating BRS computation of PARC. Note, taking advantage of fixed mode sequences is well-established in robotics and controls (e.g., [thomas2006robust, desimini2020robust, kerrigan2002optimal, wensing2023optimization]), but we believe we are the first to show its utility in the safe reach-avoid computation of robot planning models.
@@
As will be shown in Section IV-C, the use of linear interpolation to bridge between discrete and continuous time is necessary to enable avoid set computation with H-polytopes. Similar reachability techniques to ours have been shown to work with Bernstein polynomials [michaux2023can], but only for forward reachability, to the best of our knowledge. Nonetheless, it may still be possible for PARC to handle alternative interpolation methods such as cubic splines or Bézier curves by using other set representations such as polynomial zonotopes [kochdumper2020sparse]. In this paper, we opted for the most straightforward approach of linear interpolation and leave investigation of other interpolation methods as future work.
@@ II-D One Step BRS via Inverse Affine Map
Consider an affine system $\mathbf{x}(t+\Delta t)=\mathbf{C}\mathbf{x}(t)+\mathbf{d}$. Define the one-step BRS as the map $\mathcal{B}\!\left(\mathcal{P}(\mathbf{A},\mathbf{b}),\mathbf{C},\mathbf{d}\right)=\left\{\mathbf{x}|\ \mathbf{A}\mathbf{x}_{\mathrm{\textnormal{next}}}\leq\mathbf{b},\ \mathbf{x}_{\mathrm{\textnormal{next}}}=\mathbf{C}\mathbf{x}+\mathbf{d}\right\}$. This one-step BRS is in fact an H-polytope:
@@ III-A Planning Model
Changing the workspace states either does not affect the planning dynamics $\mathbf{f}_{\mathrm{\textnormal{plan}}}$ in the workspace or else the dynamics are constant:
$\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\mathbf{f}_{\mathbf{w}}\right\|_{2}=\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{p}}\right\|_{\mathrm{\textnormal{F}}}=\left\|\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{w}}\tfrac{\partial\mathbf{f}_{\mathbf{w}}}{\partial\mathbf{k}}\right\|_{\mathrm{\textnormal{F}}}=0$ (13)
@@
This is not overly restrictive; a variety of workspace-ETI planning models are shown in Section V.
@@ III-B Tracking Model
where $\mathbf{z}_{0}\in Z\subset\mathbb{R}^{{n}_{z}}$ is the initial state. The tracking state $\mathbf{z}$ includes the planning state $\mathbf{p}$, i.e., $\mathbf{z}=[\mathbf{p}^{\intercal},\mathbf{z}_{\mathrm{\textnormal{other}}}^{\intercal}]^{\intercal}$. Given a plan $\mathbf{p}$ parameterized by $\mathbf{k}$, we use a feedback tracking controller $\mathbf{u}_{\mathrm{\textnormal{fb}}}(t,\mathbf{z}(t);\mathbf{k})\in U$ (e.g., proportional–integral–derivative (PID) or model predictive control (MPC)) to control the tracking model. We assume $Z$ and $U$ are compact and the dynamics are Lipschitz such that trajectories uniquely exist [grant2014theory]. For a particular $\mathbf{z}_{0}$, we denote the resulting realized trajectory (i.e., solution to (14)) as $\mathbf{z}(t;\mathbf{z}_{0},\mathbf{k})$ at time $t$.